# All-in-one simulation-based inference

### v1

@@ Acknowledgements
: The Two Moons task is designed to test inference algorithms in handling multimodal distributions  (Greenberg et al., 2019). The prior is a Uniform distribution $U(\theta;-1,1)$. The data $x$ is generated from $\theta$ as:
$x|\theta$ $=\begin{bmatrix}r\cos(\alpha)+0.25\\ r\sin(\alpha)\end{bmatrix}+\begin{bmatrix}-|\theta_{1}+\theta_{2}|/\sqrt{2}\\ (-\theta_{1}+\theta_{2})/\sqrt{2}\end{bmatrix},$
where $\alpha\sim U(-\pi/2,\pi/2)$ and $r\sim N(0.1,0.012)$. Leading to a dimensionality $\theta\in\mathbb{R}^{2}$, $x\in\mathbb{R}^{2}$.
@@
The SLCP (Simple Likelihood Complex Posterior) task is a challenging inference task designed to generate a complex posterior distribution (Papamakarios et al., 2019; Greenberg et al., 2019; Hermans et al., 2020; Durkan et al., 2020). The setup is as follows. The prior is a uniform distribution $U(-3,3)$. The data $x$ given $\theta$ is $x|\theta=(x_{1},\ldots,x_{4})$, where each $x_{i}$ follows a Gaussian distribution $N(m_{\theta},S_{\theta})$ with:
$m_{\theta}$ $=\begin{bmatrix}\theta_{1}\\ \theta_{2}\end{bmatrix},$
$S_{\theta}$ $=\begin{bmatrix}\theta_{3}^{2}&\tanh(\theta_{5})\cdot\theta_{3}^{2}\cdot\theta_{4}^{2}\\ \tanh(\theta_{5})\cdot\theta_{3}^{2}\cdot\theta_{4}^{2}&\theta_{4}^{2}\end{bmatrix}.$
Leading to a dimensionality of $\theta\in\mathbb{R}^{5}$, $x\in\mathbb{R}^{8}$.
@@
for $i=0,...,9$. Observations are generated according to $x_{i}=\mathcal{N}(x_{i};\theta_{i}^{2},0.5**2)$, leading to a nonlinear hidden Markov model with bimodal correlated posterior and leading to a dimensionality of $\theta\in\mathbb{R}^{10}$, $x\in\mathbb{R}^{10}$.
@@
We impose a Uniform prior on the global variables, $\gamma$ and $\delta$, denoted as $\gamma,\delta\sim\text{Unif}(0,0.5)$. For the local contact rate, we adopt a stochastic approach where $\hat{\beta}$ is drawn from a sinusoidal function modulated by a Gaussian distribution, $\sin\mathcal{N}(0,k)$, with $k$ representing an RBF kernel defined as $k(t_{1},t_{2})=2.5^{2}\exp\left(-\frac{1}{2}\frac{\|t_{1}-t_{2}\|^{2}}{7^{2}}\right)$. This is further transformed via a sigmoid function to ensure $\beta(t)\in[0,1]$ for all $t$. Observational data is modeled with log-normal noise, characterized by a mean of $S(t)$ and a standard deviation of $\sigma=0.05$.
@@
where $\text{efun}(x)=\begin{cases}1-\frac{x}{2}&\text{if }x<1e-4\\ \frac{x}{\exp(x)-1.0}&\text{otherwise}\end{cases}$. This formulation leads to the comprehensive Hodgkin-Huxley differential equations:
@@
In Sec. A3.1 we include additional experiments, investigating different SDEs, compare to more methods and review efficiency. In Sec. A3.2, we review how good guidance methods can compute arbitrary conditionals, as well as general constriants.
@@
Comprehensive benchmark results have been obtained for both the Variational Exploding SDE (VESDE) and the Variance Preserving SDE (VPSDE) models, as well as for several SBI methods. These methods include Neural Posterior Estimation (NPE) (Papamakarios & Murray, 2016), Neural Likelihood Estimation (NLE) (Papamakarios et al., 2019), and Neural Ratio Estimation (NRE) (Hermans et al., 2020). The outcomes of these benchmarks are depicted in Figure A4 and Figure A5.
@@
A3.2 Investigating general guidance
@@
Diffusion guidance can vary in its implementation from less rigorous to highly rigorous approaches. Achieving rigor in this context typically necessitates a known likelihood function. However, in the realm of SBI, this likelihood function is often either intractable or exceedingly challenging to compute (Chung et al., 2023). Consequently, our focus is directed towards universally applicable approximations, as discussed in the works of Lugmayr et al. (2022) and Bansal et al. (2023).
@@
In our methodology, we integrate two principal strategies that have demonstrated efficacy in practical scenarios. The first of these strategies is self-recurrence, as advocated by Lugmayr et al. (2022). This approach has been shown to substantially improve performance, though it necessitates an increase in computational resources. The second strategy entails amalgamating the estimated score with a general constraint function, which is derived from a denoising estimate of the current state (Bansal et al., 2023; Chung et al., 2023; Rozet & Louppe, 2021). This technique exhibits remarkable flexibility in accommodating almost any arbitrary constraint. We provide a pseudo-code in Algorithm 1. In our experimental assessments it proved to be sufficiently accurate, barring scenarios involving conflicting constraints or conditions that deviate significantly from the expected distributions. For comparative purposes, we also implemented the RePaint method as proposed by Lugmayr et al. (2022). However, it is important to note that this method primarily applies to normal conditioning and does not readily extend to general constraints.
General guidance, however requires a scaling function, which up and down scales the gradient at different diffusion times $t$. As the magnitude of the marginal score does depend on the SDE, this scaling function should also. In our experiment, we generally used a scaling function of the form $s(t)=\frac{1}{\sigma(t)^{2}}$ i.e. which is inversely proportional to the variance of the approximate marginal SDE scores.

### v2

@@ Acknowledgements
: The Two Moons task is designed to test inference algorithms in handling multimodal distributions  (Greenberg et al., 2019). The prior is a Uniform distribution $U({\bm{\theta}};-1,1)$. The data ${\bm{x}}$ is generated from ${\bm{\theta}}$ as:
${\bm{x}}|{\bm{\theta}}$ $=\begin{bmatrix}r\cos(\alpha)+0.25\\ r\sin(\alpha)\end{bmatrix}+\begin{bmatrix}-|\theta_{1}+\theta_{2}|/\sqrt{2}\\ (-\theta_{1}+\theta_{2})/\sqrt{2}\end{bmatrix},$
where $\alpha\sim\mathcal{U}(-\pi/2,\pi/2)$ and $r\sim\mathcal{N}(0.1,0.012)$. Leading to a dimensionality ${\bm{\theta}}\in\mathbb{R}^{2}$, ${\bm{x}}\in\mathbb{R}^{2}$.
@@
The SLCP (Simple Likelihood Complex Posterior) task is a challenging inference task designed to generate a complex posterior distribution (Papamakarios et al., 2019; Greenberg et al., 2019; Hermans et al., 2020; Durkan et al., 2020). The setup is as follows. The prior over ${\bm{\theta}}$ is a uniform distribution $\mathcal{U}(-3,3)$. The data ${\bm{x}}$ given ${\bm{\theta}}$ is ${\bm{x}}=({\bm{x}}_{1},\ldots,{\bm{x}}_{4})$, where each ${\bm{x}}_{i}\sim\mathcal{N}(\mathbf{\mu_{\theta}},\mathbf{\Sigma_{\theta}})$ with:
$\mathbf{\mu_{\theta}}$ $=\begin{bmatrix}\theta_{1}\\ \theta_{2}\end{bmatrix},$
$\mathbf{\Sigma_{\theta}}$ $=\begin{bmatrix}\theta_{3}^{2}&\tanh(\theta_{5})\cdot\theta_{3}^{2}\cdot\theta_{4}^{2}\\ \tanh(\theta_{5})\cdot\theta_{3}^{2}\cdot\theta_{4}^{2}&\theta_{4}^{2}\end{bmatrix}.$
Leading to a dimensionality of ${\bm{\theta}}\in\mathbb{R}^{5}$, ${\bm{x}}\in\mathbb{R}^{8}$.
@@
for $i=0,...,9$. Observations are generated according to $x_{i}=\mathcal{N}(x_{i};\theta_{i}^{2},0.5^{2})$, leading to a nonlinear hidden Markov model with bimodal correlated posterior and leading to a dimensionality of ${\bm{\theta}}\in\mathbb{R}^{10}$, ${\bm{x}}\in\mathbb{R}^{10}$.
@@
We impose a Uniform prior on the global variables, $\gamma$ and $\delta$, denoted as $\gamma,\delta\sim\text{Unif}(0,0.5)$. For the time-dependent contact rate, we first sample $\hat{\beta}\sim\mathcal{G}(0,k)$ from a Gaussian process prior, with $k$ representing an RBF kernel defined as $k(t_{1},t_{2})=2.5^{2}\exp\left(-\frac{1}{2}\frac{\|t_{1}-t_{2}\|^{2}}{7^{2}}\right)$. This is further transformed via a sigmoid function to ensure $\beta(t)\in[0,1]$ for all $t$. Observational data is modeled with log-normal noise, characterized by a mean of $S(t)$ and a standard deviation of $\sigma=0.05$.
@@
where $\text{efun}(x)=\begin{cases}1-\frac{x}{2}&\text{if }x<1e-4\\ \frac{x}{\exp(x)-1.0}&\text{otherwise}\end{cases}$.
This formulation leads to the comprehensive Hodgkin-Huxley differential equations:
@@
In Sec. A3.1, we include additional experiments, i.e., investigating different SDEs, comparing to more methods, adding additional metrics, and reviewing efficiency. In Sec. A3.2, we demonstrate target inference with embedding nets on a complex task for gravitational wave data. Finally, in Sec. A3.3, we review how good guidance methods can compute arbitrary conditionals, as well as general constraints.
@@
Comprehensive benchmark results have been obtained for both the Variance Exploding SDE (VESDE) and the Variance Preserving SDE (VPSDE) models, as well as for several SBI methods. These methods include Neural Posterior Estimation (NPE) (Papamakarios & Murray, 2016), Neural Likelihood Estimation (NLE) (Papamakarios et al., 2019), and Neural Ratio Estimation (NRE) (Hermans et al., 2020). The outcomes of these benchmarks are depicted in Figure A5 and Figure A6.
@@
Average negative loglikelihood:
The average negative loglikelihood (NLL) for the true posterior is a metric suitable for evaluation on an increased number of different observations (Lueckmann et al., 2021; Hermans et al., 2022). We evaluate the average on 5000 samples from the joint distribution. We did this for both the posterior and likelihood, as estimated by Simformer, and compared it to the corresponding NPE and NLE baseline. Note that NPE and NLE are trained to minimize the NLL, giving it a natural advantage. In contrast, Simformer only indirectly minimizes negative loglikelihood through the score-matching objective. Notably, to evaluate the loglikelihood for the Simformer, we have to use the probability flow ODE (Song et al., 2021b). Hence, the loglikelihood is also based on the probability flow ODE, not the corresponding SDE formulation (which does not necessarily exactly agree for a finite number of steps). We show the corresponding result in Fig A8. In most cases, the results agree with the C2ST evaluation (which only evaluates SDE sampling quality). However, in some cases NLE or NPE does perform better with respect to this metric. The difference is due to the discrepancy between SDE sampling and ODE log probability evaluation and the fact that Simformer is not trained to minimize loglikelihood, which is not necessarily at odds with producing good samples.
Calibration:
To check whether the distributions estimated by Simformer are well-calibrated, we performed an expected coverage analysis (Hermans et al., 2022), again both for the posterior and likelihood. Intuitively, this test checks whether the ground-truth parameter lies within the top $\alpha\%$ highest density region in $\alpha\%$ of all cases (which is what the true posterior must satisfy). The same analysis was performed for NPE as a reference (see Fig. A9). In cases in which the likelihood is significantly easier to learn than the posterior (i.e., SLCP), we can observe that, indeed, the estimate of the simple likelihood becomes well calibrated earlier than the posterior (see Fig. A10, Fig. A11, Fig. A12, upper right corner). Overall, Simformer is well-calibrated and, similar to NPE, tends to more conservative approximations (coverage plots tend to be above the diagonal).
We also perform a coverage analysis on the SIR task (Fig. A13). Note that because this model is nonparametric, there are infinitely many distributions we could evaluate (i.e. by selecting different times for observations or parameters). We opt to run an evaluation on 20 random time points for each time-dependent parameter (contact rate) or observation (S, I, D).
@@
Figure A8: Average negative loglikelihood of the true parameter for NPE, NLE, and all Simformer variants. Evaluating both the likelihood (top row) and posterior (bottom row).
Figure A9: Calibration analysis for NPE using expected coverage (Hermans et al., 2022). Each row corresponds to training simulation sizes of 1k, 10k, 100k. Each column represents a task.
Figure A10: Calibration analysis for Simformer using expected coverage (Hermans et al., 2022), both for the posterior and likelihood. Each row corresponds to training simulation sizes of 1k, 10k, 100k. Each column represents a task.
Figure A11: Calibration analysis for Simformer (undirected) using expected coverage (Hermans et al., 2022), both for the posterior and likelihood. Each row corresponds to training simulation sizes of 1k, 10k, 100k. Each column represents a task.
Figure A12: Calibration analysis for Simformer (directed) using expected coverage, both for the posterior and likelihood. Each row corresponds to training simulation sizes of 1k, 10k, 100k. Each column represents a task.
Figure A13: Calibration analysis for the SIR task using expected coverage (Hermans et al., 2022). On the left, we evaluate the posterior (for randomly selected time points). On the right, we have the coverage for different conditional distributions (also for randomly selected time points).
A3.2 Targeteted inference and embedding nets
In the main manuscript, we focus on estimating all conditionals of a certain task. However, in certain scenarios, it might simply not be wanted or way harder to do so. In this case, we can query Simformer to simply target only a subset of conditionals by restricting the number of condition masks $M_{C}$ to whatever conditionals we deem worth estimating. Secondly, in tasks were data is high dimensional, it becomes computationally demanding to consider each scalar as a variable. In this case, we should encode whole vectors into a single token.
As a test case, we will consider the Gravitational Waves benchmark tasks as presented in Hermans et al. (2022). In this case, we have low dimensional $\mathbf{\theta}\in\mathbb{R}^{2}$, i.e., the masses of the two black holes, and two high dimensional ${\bm{x}}\in\mathbb{R}^{8192}$ measurements of the corresponding gravitational waves from two different detectors. In this case, it is clear that learning the likelihood, i.e., a conditional generative model for the high dimensional observations, is harder than just learning the posterior over the two parameters. A common practice for high dimensional observations is to use an embedding network, i.e., a neural network that compresses it to a lower dimensional vector. Hermans et al. (2022) did use a convolutional embedding net for NPE on this task. As already hinted in the manuscript, we can do the same for Simformer, i.e., we compress the detector measurements using a convolutional neural network into a single token. Additionally to the full posterior distribution, we are still interested in the partial posterior distributions as, e.g., there might only be measurements from one of the detectors (notably, the measurements are not independent). We hence only target the conditionals $p({\bm{\theta}}|{\bm{x}}_{1},{\bm{x}}_{2})$, $p({\bm{\theta}}|{\bm{x}}_{1})$ and $p({\bm{\theta}}|{\bm{x}}_{2})$. We use 100k simulations for training. For two examples, we show the estimated (partial) posterior(s) in Fig. A14a Fig. A14b. Simformer can combine the information from both detectors in a meaningful way (as verified by a calibration analysis, Fig. A14c.
Figure A14: Inference on gravitational wave measurements. (a) Shows the detector measurements of a gravitational wave (first column). The associated posterior and partial posteriors for the detector measurements. (b) Shows the same as in (a) but for different observations. (c) Calibration analysis of the posterior and partial posteriors in terms of expected coverage.
A3.3 Details on general guidance
@@
Diffusion guidance can vary in its implementation from less rigorous to highly rigorous approaches. Achieving rigor in this context typically necessitates a known likelihood function. However, in the realm of SBI, this likelihood function is often either intractable or challenging to compute (Chung et al., 2023). Consequently, our focus is directed towards universally applicable approximations, as discussed in the works of Lugmayr et al. (2022) and Bansal et al. (2023).
@@
In our methodology, we integrate two principal strategies that have demonstrated efficacy in practical scenarios. The first of these strategies is self-recurrence, as advocated by Lugmayr et al. (2022). This might also be interpreted as a predictor-corrector algorithm (Song et al., 2021b) with a pseudo-Gibbs sampling corrector. This approach has been shown to substantially improve performance, though it necessitates an increase in computational resources. The second strategy entails adjusting the estimated score with a general constraint function, which we evaluate on a denoised estimate of the variables (Bansal et al., 2023; Chung et al., 2023; Rozet & Louppe, 2021). Overall, this is remarkable flexibility and supports almost any constraint to be incorporated. We provide a pseudo-code in Algorithm 1. In our experimental assessments, it proved to be sufficiently accurate. For comparative purposes, we also implemented the RePaint method as proposed by Lugmayr et al. (2022). However, it is important to note that this method primarily applies to normal conditioning and does not readily extend to general constraints. On the other hand, General guidance requires the specification of a scaling function, which up and down scales the constrained score at different diffusion times $t$. As the magnitude of the marginal score does depend on the SDE, this scaling function should also. In our experiment, we generally used a scaling function of the form $s(t)=\frac{1}{\sigma(t)^{2}}$, i.e., which is inversely proportional to the variance of the approximate marginal SDE scores.